# Aspects of the chiral crossover transition in (2+1)-flavor QCD with Möbius domain-wall fermions

### v1

@@ IV.3 Comparing Möbius domain-wall results with highly-improved staggered quark data
The $\chi_{\text{top}}^{1/4}$ calculated on the gauge configurations generated with MDWF discretization on an $N_{\tau}=8$ lattice, is compared with corresponding continuum extrapolated values obtained by using the HISQ discretization from Ref. [27] in Fig. 6. The topological charge on the HISQ configurations was obtained using the improved clover definition, after performing Symanzik flow to reduce the ultra-violet noise resulting from gauge fluctuations [68]. We observe that, even though the results obtained from the MDWF ensembles are for a fixed lattice spacing, these are consistent with the continuum extrapolated results obtained using HISQ fermions. It is to be noted that the calculations involving MDWF consisted of $\mathcal{O}(10^{2})$ i.e. only about a tenth of the configurations we have generated, whereas the HISQ results were obtained using $\mathcal{O}(10^{3})$ configurations. Our results demonstrate a first comparison of the topological susceptibility, at finite temperature, of continuum extrapolated HISQ results with MDWF, highlighting the importance of improved realization of chiral symmetry on the lattice also for a proper topological sampling among the gauge configurations.
V Summary & Outlook
@@
In this paper we have shown the importance of the MDWF discretization, which respects the chiral symmetry to a good extent on a finite lattice, for understanding the thermodynamics of the chiral crossover transition in QCD. Since the light quark mass is significantly smaller than any other scale in finite temperature QCD describing physical quarks, it is sensible to revisit the question whether the singlet and non-singlet sub-groups of chiral symmetry are effectively restored simultaneously. Addressing this question on the lattice using the HISQ discretization is more subtle, since it breaks the non-singlet chiral symmetry group to a U(1) subgroup, as a result of a mixing between the spin-flavor degrees of freedom. Further, due to lack of an exact index of the HISQ Dirac matrix, the topological fluctuations, which are responsible for the effective restoration of the singlet part of the chiral symmetry in the continuum, cannot be explicitly related to each other on the lattice. While taking the continuum limit of chiral observables such as $\chi_{\text{disc}}$ calculated using the HISQ discretization is strictly necessary to extract the corresponding pseudo-critical temperature, we have shown that the MWDF discretization reproduces the continuum QCD results well already on a finite $N_{\tau}=8$ lattice. It is therefore expected that calculating $\chi_{\text{disc}}$ using MDWF on a finer lattice, would show only a weak cutoff dependence, since the chiral and continuum limits are very well disentangled, in contrast to staggered fermions. Furthermore, we find that the HISQ results for $\chi_{\text{disc}}$ on an $N_{\tau}=8$ lattice have large cutoff effects and cannot be taken into consideration in the process of estimating its continuum values using an $\mathcal{O}(a^{2})$ extrapolation. We have also shown that the topological fluctuations quantified by the observable $\chi_{\text{top}}$, obtained from calculations on the $N_{\tau}=8$ MDWF gauge ensembles, are similar in magnitude to the continuum extrapolated values calculated on the HISQ gauge ensembles. This highlights once more the role of chiral symmetry in also realizing the topological fluctuations more accurately.
@@
By determining the exponent $\alpha$ controlling the temperature dependence of $\chi_{\text{top}}$ using MDWF calculations, we have shown that at $T\approx 185\,\mathrm{MeV}$ the singlet part of the chiral symmetry is not yet effectively restored, i.e. to its asymptotically high temperature value given by a dilute instanton gas approximation. This finding suggests that the chiral phase transition in the massless limit of (2+1)-flavor QCD is of second order with critical exponents belonging to the $\mathrm{O}(4)$ universality class.
@@
Apart from extracting the pseudo-critical temperature $T_{pc}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$ from the peak of the disconnected chiral susceptibility $\chi_{\text{disc}}$, we independently show that the non-singlet part of chiral symmetry is also effectively restored in the same temperature range. This is evident from the deviation of $\chi_{\text{top}}^{1/4}/T_{pc}$ from the chiral perturbation theory prediction and from the realization of the relation $\chi_{\text{top}}=m_{\ell}^{2}\chi_{\text{disc}}$ up to $\mathcal{O}(m_{\ell}^{4})$ corrections due to the finite mass of the light quarks.
@@
While it is anticipated that the cutoff dependence of chiral observables, such as $\chi_{\mathrm{disc}}$, computed using the MDWF discretization is much weaker compared to calculations using the HISQ action, which suffer from taste violation effects, it would nonetheless be instructive to explicitly verify this hypothesis by performing our computations and analysis on a finer lattice. It would also be interesting to directly calculate the scalar iso-triplet susceptibility through precise computation of the connected chiral susceptibility. This would require additional inversions of the Dirac matrix. Furthermore, it would be important to study in detail the properties of the eigenvalues and eigenvectors [52, 69, 28, 70] of the MDWF operator in order to better understand the mechanism underlying the chiral symmetry restoration from the temperature dependence of the eigenvalue density.
Acknowledgements
@@
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Proj. No. 315477589-TRR 211; the PUNCH4NFDI consortium supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) with project number 460248186 (PUNCH4NFDI). MS acknowledges support by the Taiwanese NSTC grant 113-2639-M-002-006-ASP. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Additional computations have been performed on the GPU Cluster at Bielefeld University and at the Institute of Mathematical Sciences, Chennai. Our codes were developed in part using the Grid Python Toolkit (GPT) [71] which uses the Grid library [72]. Part of the data analysis has been performed using routines from the AnalysisToolbox [73] code developed by the HotQCD Collaboration. Part of the HISQ calculations were performed using the SIMULATeQCD [68, 74] library. We also thank Christoph Lehner for his advice and help in using GPT, and Yu Zhang for helpful discussions on DWF calculations.

### v2

@@ IV.3 Comparing chiral observables in the light quark sector calculated with MDWF and HISQ
The $\chi_{\text{top}}^{1/4}$ calculated on the gauge configurations generated with MDWF discretization on an $N_{\tau}=8$ lattice, is compared with corresponding continuum extrapolated values obtained by using the HISQ discretization from Ref. [29] in Fig. 6. The topological charge on the HISQ configurations was obtained using the improved clover definition, after performing Symanzik flow to reduce the ultra-violet noise resulting from gauge fluctuations [71]. We observe that, even though the results obtained from the MDWF ensembles are for a fixed lattice spacing, these are consistent with the continuum extrapolated results obtained using HISQ fermions. It is to be noted that the calculations involving MDWF consisted of $\mathcal{O}(10^{2})$ i.e., only about a tenth of the configurations we have generated, whereas the HISQ results were obtained using $\mathcal{O}(10^{3})$ configurations. Our results demonstrate a first comparison of the topological susceptibility, at finite temperature, of continuum extrapolated HISQ results with MDWF, highlighting the importance of improved realization of chiral symmetry on the lattice also for a proper topological sampling among the gauge configurations.
V Summary and Outlook
@@
In this paper we have shown the importance of the MDWF discretization, which respects the chiral symmetry to a good extent on a finite lattice, for understanding the thermodynamics of the chiral crossover transition in QCD. Since the light quark mass is significantly smaller than any other scale in finite temperature QCD describing physical quarks, it is sensible to revisit the question whether the singlet and nonsinglet subgroups of chiral symmetry are effectively restored simultaneously. Addressing this question on the lattice using the HISQ discretization is more subtle, since it breaks the nonsinglet chiral symmetry group to a U(1) subgroup, as a result of a mixing between the spin-flavor degrees of freedom. Further, due to lack of an exact index of the HISQ Dirac matrix, the topological fluctuations, which are responsible for the effective restoration of the singlet part of the chiral symmetry in the continuum, cannot be explicitly related to each other on the lattice. While taking the continuum limit of chiral observables such as $\chi_{\text{disc}}$ calculated using the HISQ discretization is strictly necessary to extract the corresponding pseudocritical temperature, we have shown that the MDWF discretization reproduces the continuum QCD results well already on a finite $N_{\tau}=8$ lattice. It is therefore expected that calculating $\chi_{\text{disc}}$ using MDWF on a finer lattice, would show only a weak cutoff dependence, since the chiral and continuum limits are very well disentangled, in contrast to staggered fermions. Furthermore, we find that the HISQ results for $\chi_{\text{disc}}$ on an $N_{\tau}=8$ lattice have large cutoff effects and cannot be taken into consideration in the process of estimating its continuum values using an $\mathcal{O}(a^{2})$ extrapolation. We have also shown that the topological fluctuations quantified by the observable $\chi_{\text{top}}$, obtained from calculations on the $N_{\tau}=8$ MDWF gauge ensembles, are similar in magnitude to the continuum extrapolated values calculated on the HISQ gauge ensembles. This highlights once more the role of chiral symmetry in also realizing the topological fluctuations more accurately.
@@
By determining the exponent $\alpha$ controlling the temperature dependence of $\chi_{\text{top}}$ using MDWF calculations, we have shown that at $T\approx 185\,\mathrm{MeV}$ the singlet part of the chiral symmetry is not yet effectively restored, i.e., to its asymptotically high temperature value given by a dilute instanton gas approximation. This finding suggests that the chiral phase transition in the massless limit of (2+1)-flavor QCD is of second order with critical exponents belonging to the $\mathrm{O}(4)$ universality class.
@@
Apart from extracting the pseudocritical temperature $T_{pc}=158.7{}_{{}\mathop{-}2.3}^{{}\mathop{+}2.6}\,\mathrm{MeV}$ from the peak of the disconnected chiral susceptibility $\chi_{\text{disc}}$, we independently show that the nonsinglet part of chiral symmetry is also effectively restored in the same temperature range. This is evident from the deviation of $\chi_{\text{top}}^{1/4}/T_{pc}$ from the chiral perturbation theory prediction and from the realization of the relation $\chi_{\text{top}}=m_{\ell}^{2}\chi_{\text{disc}}$ up to $\mathcal{O}(m_{\ell}^{4})$ corrections due to the finite mass of the light quarks.
@@
While it is anticipated that the cutoff dependence of chiral observables, such as $\chi_{\mathrm{disc}}$, computed using the MDWF discretization is much weaker compared to calculations using the HISQ action, which suffer from taste violation effects, it would nonetheless be instructive to explicitly verify this hypothesis by performing our computations and analysis on a finer lattice. It would also be interesting to directly calculate the scalar isotriplet susceptibility through precise computation of the connected chiral susceptibility. This would require additional inversions of the Dirac matrix. Furthermore, it would be important to study in detail the properties of the eigenvalues and eigenvectors [55, 72, 30, 73] of the MDWF operator in order to better understand the mechanism underlying the chiral symmetry restoration from the temperature dependence of the eigenvalue density.
Acknowledgments
@@
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Proj. No. 315477589-TRR 211; the PUNCH4NFDI consortium supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) with project number 460248186 (PUNCH4NFDI). M..S acknowledges support by the Taiwanese NSTC Grant No. 113-2639-M-002-006-ASP. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. [74] for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Additional computations have been performed on the GPU Cluster at Bielefeld University and at the Institute of Mathematical Sciences, Chennai. Our codes were developed in part using the Grid Python Toolkit (GPT) [75] which uses the Grid library [76]. Part of the data analysis has been performed using routines from the AnalysisToolbox [77] code developed by the HotQCD Collaboration. Part of the HISQ calculations were performed using the SIMULATeQCD [71, 78] library. We also thank Christoph Lehner for his advice and help in using GPT, and Yu Zhang for helpful discussions on DWF calculations.
Data Availability
The data that support the findings of this work are openly available [79].